There are 7 notes of "A" on a standard piano. When the third A is struck, the string associated with the key vibrates 220 times per second. The next A above the given A vibrates twice as fast. An exponential function with a base of two is used to determine the frequency of the 11 notes between the two A's.

Find this function. Hint: The function is of the form f(x) = k · 2(cx),where k and c are constants. Also f(0) = 220 and f(12) = 440.
1
Expert's answer
2020-09-23T18:21:49-0400

The function is of the form "f(x) = k" "\u00b72^{cx}", so we need to find "k" and "c"

Also we know "f(0) = 220, f(12) = 440"

Create the system of equations:


"220 = k \u00b7 2^{c*0}"

"440 = k \u00b7 2^{c*12}"


From first equations: "k = 220"

From second equations:

"440 = 220 * 2^{12c}"

"2 = 2^{12c}"

"12c = 1"

"c=\frac{1}{12}"


So our function is: "f(x) = 220 * 2^{\frac{1}{12}x}"